User Tools

Site Tools



This shows you the differences between two versions of the page.

Link to this comparison view

Both sides previous revision Previous revision
Next revision
Previous revision
lbaops:lbajul2011:v252yholog [2011/07/21 07:28]
lbaops:lbajul2011:v252yholog [2011/07/22 15:20]
Line 13: Line 13:
 Frequency counter reading 0.01702 ​ Frequency counter reading 0.01702 ​
-fmout-gps = +2.04067e-05+fmout-gps = +20.4 microseconds
 John Dickey observer John Dickey observer
 +At around 2330, Noise diode problem was fixed. The noise diode control had not been set to remote before this and so Tsys measurements in the FS log would not have been working.
 +Brett Reid.
 +Stopped recording about 5 min early, no other problems.
 +(fringes found).
 +Claire Hotan
lbaops/lbajul2011/v252yholog.txt · Last modified: 2015/12/18 16:38 (external edit)